Loan - EUR 2,500
issued for 6 years
compound interest rate of 7.76%,
What is the amount of the periodic payment which is made at the beginning of the year?

If the money were paid at the end of each year, what would be the result?
1
Expert's answer
2020-12-30T03:12:40-0500

PV=2500

at the beginning of the year:

"PV=C[\frac{1-(1+r)^{-n}}{r}]\times(1+r)"

"2500=C[\frac{1-(1+0.0776)^{-6}}{0.0776}]\times(1+0.0776)"

"C=\frac{2500}{5.01806503}=498.2"


at the end of each year:

"PV=C[\frac{1-(1+r)^{-n}}{r}]"

"2500=C[\frac{1-(1+0.0776)^{-6}}{0.0776}]"

"C=\frac{2500}{4.65673159}=536.86"